# Towards Evaluating Large Language Models on Sarcasm Understanding

### v1

@@ 4.5 Impact of Number of Demonstrations
In few-shot settings, we investigate the impact of the number of demonstrations for the state-of-the-art LLM, namely whether GPT-4 could perform better or not if more contextual examples are provided. We design five $k$-shot settings: zero-shot, two-shot, four-shot, six-shot, eight-shot. For each setting, we selectly sample $k=\left\{0,2,4,6,8\right\}$ examples that are similar to the test sample using kNN sampling algorithm. The results are shown in Fig. 4.
@@
As we can see, the number of demonstrations (k-shot) has a significant impact on the F1 scores for both Mistral and GPT-4 in the sarcasm classification task. For instance, Mistral shows an overall upward trend in F1 scores from 0-shot to 8-shot, increasing from 51.6 to 56.2, representing a total improvement of 8.7%. The most significant performance boost occurs at 2-shot, reaching 55.8, which is an 8.2% increase compared to 0-shot. This shows that Mistral has greater sensitivity to the number of examples provided, benefiting more from additional examples.
@@
In contrast, GPT-4 consistently outperforms Mistral across all shot numbers, with a baseline performance of 64.5 F1 score. GPT-4 achieves its best performance at 2-shot, with an F1 score of 67.5, showing a 4.6% improvement over 0-shot. However, it it experiences a slight decline in performance at 6-shot (63.4) and 8-shot (64.3). This indicates that GPT-4 may reach its peak performance with fewer examples and does not benefit as much from additional demonstrations.
@@
These results suggest that both models reach their optimal performance with 2-4 samples, indicating that moderate few-shot learning may be more effective than using a large number of samples for sarcasm detection tasks.
4.6 Chain of Thought Results
@@
In few-shot settings, we investigate a widely used prompting approach, CoT. The experimental results are shown in Table 4 and Fig. 6. Despite that CoT prompting has significant advantage over standard IO prompting across logical, mathematical, and commonsense reasoning tasks, it may be not sufficient in solving sarcasm understanding tasks. For example, GPT-4 sees a decline in its average F1 score from 68.4 with IO prompting to 64.7 with CoT prompting, a 5.5% decrease. Similarly, Mistral, which benefits from IO prompting with an F1 score of 55.8, drops to 47.5 with CoT prompting, reflecting a 14.88% reduction. These examples highlight that while CoT can improve reasoning, it may introduce complexity that hinders models’ ability to effectively detect and interpret sarcasm. Qwen 2 sees a 25.1% reduction, from 53.1 to 39.7. These results indicate that CoT often complicates sarcasm understanding, negatively impacting performance. We present an intuitive comparison in Fig. 7.
@@
The reason is that sarcasm detection is often considered a holistic and non-rational cognitive process that does not conform to step-by-step logical reasoning. Sarcasm expression does not strictly conform to formal logical structures, such as the law of hypothetical syllogism (i.e., if $\mathcal{A}\Rightarrow\mathcal{B}$ and $\mathcal{B}\Rightarrow\mathcal{C}$, then $\mathcal{A}\Rightarrow\mathcal{C}$). For example, “Poor Alice has fallen for that stupid Bob; and that stupid Bob is head over heels for Claire; but don’t assume for a second that Alice would like Claire”. Hence, CoT prompting is less effective for sarcasm classification compared to the more straightforward IO approach. There is an urgent need to develop sarcasm understanding prompting approaches.
@@ 4.7 Error Analysis
The detailed error analysis is also conducted via the confusion matrices that are shown in Fig. 8, Fig. 9 and Fig. 10. Each cell $\left(i,j\right)$ represents the percentage of class $i$ is classified to be class $j$. Upon reviewing the classification results produced by GPT-4 Turbo on six datasets with three prompting approaches, we discover that imbalanced categories and the sarcastic samples are the key factors contributing to misclassification.
In zero-shot IO setting, GPT-4 Turbo performs well in identifying non-sarcastic instances, particularly on datasets like IAC-V2 and SemEval Task 3, where the true positive rates for non-sarcastic labels are very high. However, the model struggles more with distinguishing sarcastic instances, especially on datasets like IAC-V1 and iSarcasmEval, where the true positive rates for sarcastic labels are lower (0.46 and 0.38, respectively). This indicates that without any prior examples, the model tends to lean towards non-sarcastic predictions, failing to capture the nuanced language that often accompanies sarcasm.
In few-shot IO setting, we observe the improvement in performance across the datasets. The model benefits from a few examples, particularly in non-sarcastic classifications. In datasets like IAC-V2, the model still struggles, with only 46% of sarcastic instances correctly identified. The model shows improvement but remains inconsistent in sarcastic classifications, indicating that while few-shot prompting provides some benefit, it may not be sufficient to fully capture sarcasm’s complexity.
In few-shot CoT setting, GPT-4 Turbo also similar difficulties in distinguishing between sarcastic and non-sarcastic instances, particularly in datasets with more subtle or nuanced sarcastic content.The reason is that CoT is not suitable to solve human sarcasm understanding tasks.
@@ 4.9 Evaluating LLMs on Multi-Modal Sarcasm Understanding
Due to the rapid surge of multimodal data on social networks, leveraging multi-modal information to enhance human language understanding and reasoning has become increasingly attractive and significant. Beyond text-based LLMs, multi-modal LLMs typically incorporate modality encoders, connectors, and generators. Through modality-aligned pre-training, these models acquire the ability to process multimodal information. Therefore, in this section, we aim to investigate their capabilities for understanding multi-modal sarcasm.
More specifically, we select one popular and publicly accessible multi-modal sarcasm detection dataset, namely, MMSD [48], as our experimental platform. We evaluate one state-of-the-art multi-modal LLM: LLaVA 1.5 [49]. Given that CoT prompting is less effective in addressing sarcasm understanding tasks, we employ only the zero-shot IO prompting method. Additionally, we present four state-of-the-art multi-modal sarcasm detection baselines for comparison. The results are shown in Table 6.
@@
MMSD
Paradigm Model Acc P R F1
MMMA [16] 86.7 80.2 84.4 82.6
ResNet+BERT [51] 84.8 77.8 84.1 80.8
MILNet [52] 89.5 85.2 89.2 87.1
Multi-Modal PLMs Multi-view CLIP [53] 88.3 82.7 88.7 85.6
LLaVA 1.5 93.7 93.7 93.1 93.4
We can notice that LLaVA 1.5 significantly outperforms other models in multi-modal sarcasm detection across all metrics—accuracy, precision, recall, and F1 score. This superior performance underscores LLaVA 1.5’s enhanced capability to integrate and interpret visual and textual cues effectively. The comparative analysis reveals that while traditional multi-modal models are effective, the advanced training methodologies and deeper contextual understanding in LLaVA 1.5 provide a more refined analysis of sarcasm.
@@ 5 Conclusion
In the era of large language models (LLMs), there is growing concern that LLMs’ success may not be fully tenable when considering sarcasm understanding. To address this question, we select eleven SOTA LLMs and eight SOTA pre-trained language models (PLMs) and present comprehensive evaluations on six widely used benchmark datasets. The results show that current LLMs underperform supervised PLMs based sarcasm detection baselines across six sarcasm benchmarks. In addition, GPT-4 consistently and significantly outperforms other LLMs across various prompting methods. This suggests that significant efforts are still required to improve LLMs’ understanding of human sarcasm. In the future, we plan to improve the classification performance by integrating the predictions of multiple models using methods such as bagging or boosting.

### v2

@@ 4.7 Impact of Number of Demonstrations
In few-shot settings, we investigate the impact of the number of demonstrations for the state-of-the-art LLM, namely whether GPT-4 could perform better or not if more contextual examples are provided. We design five $k$-shot settings: zero-shot, two-shot, four-shot, six-shot, eight-shot. For each setting, we select sample $k=\left\{0,2,4,6,8\right\}$ examples that are similar to the test sample using the kNN sampling algorithm. The results are shown in Fig. 4.
@@
As we can see, the number of demonstrations (k-shot) has a significant impact on the F1 scores for both Mistral and GPT-4 in the sarcasm classification task. For instance, Mistral shows an overall upward trend in F1 scores from 0-shot to 8-shot, increasing from 51.6 to 56.2, representing a total improvement of 8.7% $\uparrow$. The most significant performance boost occurs at 2-shot, reaching 55.8, which is an 8.2% increase compared to 0-shot. This shows that Mistral has greater sensitivity to the number of examples provided, benefiting more from additional examples.
@@
In contrast, GPT-4 consistently outperforms Mistral across all shot numbers, with a baseline performance of 64.5 F1 score. GPT-4 achieves its best performance at 2-shot, with an F1 score of 67.5, showing a 4.6% improvement over 0-shot. However, it experiences a slight decline in performance at 6-shot (63.4) and 8-shot (64.3). This indicates that GPT-4 may reach its peak performance with fewer examples and does not benefit as much from additional demonstrations.
@@
These results suggest that both models reach their optimal performance with 2-4 samples, indicating that moderate few-shot learning may be more effective than using a large number of samples for sarcasm detection tasks. Sarcasm detection can vary significantly across textual contexts and datasets, potentially explaining the variations in optimal demonstration numbers.
4.8 CoT Results
@@
In few-shot settings, we investigate a widely used prompting approach, CoT. The experimental results are shown in Table 4 and Fig. 5. Despite that CoT prompting has a significant advantage over standard IO prompting across logical, mathematical, and commonsense reasoning tasks, it may be not sufficient in solving sarcasm understanding tasks. For example, GPT-4 sees a decline in its average F1 score from 68.4 with IO prompting to 64.7 with CoT prompting, a 5.5% decrease. Similarly, Mistral, which benefits from IO prompting with an F1 score of 55.8, drops to 47.5 with CoT prompting, reflecting a 14.88% reduction. These examples highlight that while CoT can improve reasoning, it may introduce complexity that hinders models’ ability to effectively detect and interpret sarcasm. Qwen 2 sees a 25.1% reduction, from 53.1 to 39.7. These results indicate that CoT often complicates sarcasm understanding, negatively impacting performance. We present an intuitive comparison in Fig. 6.
@@
The reason is that sarcasm detection is often considered a holistic and non-rational cognitive process that does not conform to step-by-step logical reasoning. Sarcasm expression does not strictly conform to formal logical structures, such as the law of hypothetical syllogism (i.e., if $\mathcal{A}\Rightarrow\mathcal{B}$ and $\mathcal{B}\Rightarrow\mathcal{C}$, then $\mathcal{A}\Rightarrow\mathcal{C}$). For example, “Poor Alice has fallen for that stupid Bob; and that stupid Bob is head over heels for Claire; but don’t assume for a second that Alice would like Claire”. Another possible reason is that CoT does not explicitly harnesses the quintessential property of sarcasm (namely the contradiction between surface sentiment and true intention). Hence, CoT prompting is less effective for sarcasm classification compared to the more straightforward IO approach. There is an urgent need to develop sarcasm understanding prompting approaches.
@@ 4.9 Error Analysis
The detailed error analysis is also conducted via the confusion matrices that are shown in Fig. 7, Fig. 8 and Fig. 9. Each cell $\left(i,j\right)$ represents the percentage of class $i$ that is classified to be class $j$. Upon reviewing the classification results produced by GPT-4 Turbo on six datasets with three prompting approaches, we discover that imbalanced categories and the sarcastic samples are the key factors contributing to misclassification.
In the zero-shot IO setting, GPT-4 Turbo performs well in identifying non-sarcastic instances, particularly on datasets like IAC-V2 and SemEval Task 3, where the true positive rates for non-sarcastic labels are very high. However, the model struggles more with distinguishing sarcastic instances, especially on datasets like IAC-V1 and iSarcasmEval, where the true positive rates for sarcastic labels are lower (0.46 and 0.38, respectively). This indicates that without any prior examples, the model tends to lean towards non-sarcastic predictions, failing to capture the nuanced language that often accompanies sarcasm.
In the few-shot IO setting, we observe the improvement in performance across the datasets. The model benefits from a few examples, particularly in non-sarcastic classifications. In datasets like IAC-V2, the model still struggles, with only 46% of sarcastic instances correctly identified. The model shows improvement but remains inconsistent in sarcastic classifications, indicating that while few-shot prompting provides some benefit, it may not be sufficient to fully capture sarcasm’s complexity.
In the few-shot CoT setting, GPT-4 Turbo also similar difficulties in distinguishing between sarcastic and non-sarcastic instances, particularly in datasets with more subtle or nuanced sarcastic content.The reason is that CoT is not suitable to solve human sarcasm understanding tasks.
@@ 4.11 Evaluating LLMs on Multi-Modal Sarcasm Understanding
Due to the rapid surge of multi-modal data on social networks, leveraging multi-modal information to enhance human language understanding and reasoning has become increasingly attractive and significant. Beyond text-based LLMs, multi-modal LLMs typically incorporate modality encoders, connectors, and generators. Through modality-aligned pre-training, these models acquire the ability to process multi-modal information. Therefore, in this section, we aim to investigate their capabilities for understanding multi-modal sarcasm.
More specifically, we select two popular and publicly accessible multi-modal sarcasm detection datasets, namely, MMSD [46] and CMMA [17], as our experimental platforms. We evaluate four state-of-the-art multi-modal LLMs: Retrieval-LLaVA 1.5 [47], GPT-4V [^7] https://openai.com/index/gpt-4v-system-card/, Wenxin 4888https://yiyan.baidu.com/, and Qwen-VL-Plus [48]. Given that CoT prompting is less effective in sarcasm understanding, we employ only the zero-shot IO prompting method for GPT-4V, Qwen-VL-Plus and Wenxin 4. Their prompts are: “Based on the image and corresponding text below, analysis both the image and the text, determine whether they are sarcastic or not. If they are sarcastic output ‘sarcastic’, otherwise, output ‘non-sarcastic’. Return the label only without any other text.” Retrieval-LLaVA 1.5 proposes a retrieval module for LLaVA 1.5 to search for demonstrations, aiming at further bridging the gap between LLaVA 1.5 and the specific multi-modal sarcasm detection task. Additionally, we present four state-of-the-art multi-modal sarcasm detection baselines for comparison. The results are shown in Table 6.
We can notice: (1) Zero-shot IO prompting-based multi-modal LLMs exhibit weak performance across both datasets when compared to PLMs. This suggests a possible limitation in their generalizability or adaptability to complex multi-modal tasks without additional contextual support. (2) Among the LLMs tested, GPT-4V shows superior results on the MMSD dataset, which is English, compared to Wenxin 4 and Qwen-VL-Plus. In contrast, Qwen-VL-Plus excels on the CMMA dataset, which is Chinese. This divergence likely stems from the inherent differences in how sarcasm is expressed and contextualized in English versus Chinese. This indicates that GPT-4V may be more suitable for processing sarcasm in English while Qwen-VL-Plus may be suited for Chinese sarcasm understanding. (3) The superior performance of Retrieval-LLaVA 1.5, particularly notable when it utilizes a retrieval approach to incorporate relevant examples into its prompts. This model outperforms the four PLMs significantly, demonstrating that adding contextually appropriate examples can substantially enhance model performance. We also present several examples in Fig. 7.
@@
MMSD CMMA
Paradigm Model Acc P R F1 Acc P R F1
MMMA [17] 86.7 80.2 84.4 82.6 77.5 76.3 74.2 75.2
ResNet+BERT [49] 84.8 77.8 84.1 80.8 69.4 67.8 65.7 66.8
MILNet [50] 89.5 85.2 89.2 87.1 - - - -
Multi-Modal PLMs Multi-view CLIP [51] 88.3 82.7 88.7 85.6 - - - -
Retrieval-LLaVA 1.5 [47] 90.0 89.3 89.6 89.4 - - - -
Wenxin 4 60.0 52.6 31.5 39.4 50.0 40.0 50.0 44.4
Qwen-VL-Plus 64.4 52.7 95.0 67.8 39.0 39.4 97.5 56.1
Multi-Modal LLMs GPT-4V 71.0 59.4 96.5 73.5 52.0 31.8 17.6 22.6
Table 7: Typical multi-modal examples for case study.
Example Image Text Golden GPT-4V Qwen-VL WenXin4
1 [Uncaptioned image] happy new year , everyone ! xoxo # year2016 # partyhard # nihilistmemes sarcastic ×
2 [Uncaptioned image] the pa welcome center is hopping today . sarcastic ×
3 [Uncaptioned image] feeding my abs nothing but the best quality beef sarcastic ×
4 [Uncaptioned image] because every mildly successful cgi film needs an animated spinoff . sarcastic ×
5 [Uncaptioned image] when ryan seacrest air-kisses you , but you went for the real kiss non-sarcastic × × ×
6 [Uncaptioned image] bae and i …… all day . non-sarcastic
@@ 5 Conclusion
In the era of large language models (LLMs), there is growing concern that LLMs’ success may not be fully tenable when considering sarcasm understanding. To address this question, we select eleven SOTA LLMs and eight SOTA PLMs and present comprehensive evaluations on six widely used benchmark datasets. The results show that current LLMs underperform supervised PLMs based sarcasm detection baselines across six sarcasm benchmarks. In addition, GPT-4 consistently and significantly outperforms other LLMs across various prompting methods. This suggests that significant efforts are still required to improve LLMs’ understanding of human sarcasm. In the future, we plan to improve the classification performance by integrating the predictions of multiple models using methods such as bagging or boosting.